Is Nathan and Shane's age together is 58 and Nathan's age is 8 less than half of Shane's age what are both here ages
Answer by Cromlix(4381)

Hi there,
Is Nathan and Shane's age together is 58
N + S = 58
Nathan's age is 8 less than half of Shane's age
N = S/2 - 8
Using equation:
N + S = 58
Substitute
S/2 - 8 for N
S/2 - 8 + S = 58
Multiply through by 2
S - 16 + 2S = 116
Collect like terms
S + 2S = 116 + 16
3S = 132
S = 44
Nathan = 58 - 44 = 14 yrs old
Shane = 44 yrs old.
Hope this helps:-)